(i)    𝑦= 2𝑥2 + 12𝑥 + 18



1
Expert's answer
2022-03-15T18:07:00-0400

Let y = 0

2x² +12x + 18= 0

0 = 2x² + 12x + 18


0 = x² + 6x +9


0 = x² + 3x + 3x +9


0 = x(x+3) + 3(x+3)


0 = (x+3)(x+3)

x+3 =0

x = -3

Therefore, y= 0, x = -3